OpenStudy (anonymous):

Solving Quadratic Equations By Factoring. (x-1)(x+7)=0 x=?

OpenStudy (anonymous):

it is already factored

OpenStudy (anonymous):

here is a good way of looking at it A*B=0 (x-1)*B=0 x-1=0 would make 0*B=0 no matter what B holds Answer x=1 And A*(x+7) =0 x+7=0 would be A*0=0 no matter what A holds x=-7 as you can see solutions are 1 and -7

OpenStudy (anonymous):

oh yea that does make sense, thanks

OpenStudy (anonymous):

OH in general (x-a)(x-b)(x-c)(x-d)... and so on a, b, c, d... are the solutions. The trick is to get it that way tho :P

OpenStudy (anonymous):

thanks for the trick, but i would have too answer it like this (x+1) and (x-7). is it saying we basically answer the positive to negative and negative to positive?

OpenStudy (anonymous):

You could look at it like that. Here are examples x+a=0 -a would be the answer x-a =0 a would be the answer x-1=0 1 would be the answer x+1=0 -1 would be the answer

OpenStudy (anonymous):

yes, because x is also 1

OpenStudy (anonymous):

Yup